# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

### v1

@@ Abstract
Autonomous driving has rapidly developed and shown promising performance with recent advances in hardware and deep learning methods. High-quality datasets are fundamental for developing reliable autonomous driving algorithms. Previous dataset surveys tried to review the datasets but either focused on a limited number or lacked detailed investigation of the characters of datasets. To this end, we present an exhaustive study of over 200 autonomous driving datasets from multiple perspectives, including sensor modalities, data size, tasks, and contextual conditions. We introduce a novel metric to evaluate the impact of each dataset, which can also be a guide for establishing new datasets. We further analyze the annotation process and quality of datasets. Additionally, we conduct an in-depth analysis of the data distribution of several vital datasets. Finally, we discuss the development trend of the future autonomous driving datasets.
@@
Dataset, influence, annotation, autonomous driving.
@@ I Introduction
Autonomous driving (AD) aims to revolutionize the transportation system by creating vehicles that can accurately perceive their environment, make intelligent decisions, and drive safely without human intervention. Due to thrilling technical development, various autonomous driving products have been implemented in several fields, such as robotaxi [1]. These rapid advancements in autonomous driving rely heavily on extensive datasets, which help autonomous driving systems be robust and reliable in complex driving environments.
In recent years, there has been a significant increase in the quality and variety of autonomous driving datasets. The first apparent phenomenon in the development of datasets is the various data collection strategies, including synthetic datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] generated by simulators and recorded from the real world [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], to name just a few. Secondly, the datasets vary in composition, including but not limited to multiple sensory data (like camera images and LiDAR point clouds) and different annotation types for various tasks in autonomous driving. Fig. 1 depicts the statistic of the 3D object bounding box distribution of six real-world datasets (Argoverse 2[28], KITTI [13], nuScenes [22], ONCE [29], Waymo [23], and ZOD [30]) under a bird’s-eye-view perspective, showing each dataset’s distinguished annotation characters. According to the equipment position of sensors, the variety of datasets also reflects in the sensing domains, containing onboard, V2X, drone, and others. Furthermore, the geometrical diversity and weather conditions improve the generalization of the autonomous driving datasets.
Figure 1: BEV object distribution of datasets. Each heatmap corresponds to a dataset plotted with X and Y coordinates. Y is the driving direction of the ego-vehicle. The unique annotation characters of each dataset are reflected in the distribution range, density, and number of bounding boxes.
@@ I-A Research Gap & Motivation
We demonstrate the yearly published number of perception datasets in Fig. 2 to reflect the trend of autonomous driving datasets from one perspective. Since numerous and continually increasing numbers of publicly published datasets exist, an exhaustive survey on autonomous driving datasets is valuable for advancing academic and industrial research. In prior work, Yin et al. [31] summarized 27 publicly available datasets containing data collected on public roads. As a sequential work of  [31], [32] extended the dataset number. [33] and [34] proposed systematical introductions to the existing datasets from an application perspective. Beyond describing existing datasets, [35] discussed the domain adaptation between synthetic and real data and automatic labeling methods. [36] summarized existing datasets and undertook an exhaustive analysis of the characters of the next-generation datasets. However, these surveys only summarized a small number of datasets, causing a non-wide scope. AD-Dataset [37] collected a large number of datasets while lacking detailed analysis for the attributes of these datasets. Compared to studies on all types of datasets, some researchers presented surveys on a particular type of autonomous driving dataset, such as anomaly detection [38], synthetic datasets [39], 3D semantic segmentation [40], and decision-making [41].
To this end, we aim to propose a comprehensive and systematic study on a large number of datasets in autonomous driving, covering all tasks from perception to control, considering real-world as well as synthetic data, and providing insight into the data modality and quality of several crucial datasets. We illustrate the comparison between other dataset surveys and ours in Tab I.
@@ A Survey on AD Datasets [35] 2021 30 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ours 235 Perc, Pred, Pl, C, E2E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TABLE I: We compare our survey paper with other AD dataset surveys in the following perspectives: collected dataset number (#Dataset), relevant tasks, sensing domain (S. domain), sensor modality (S. moda.), geometric conditions (Geo.), environmental conditions (Env.), analyzing data distribution, introducing annotation quality and process. We describe the task types in a coarse granularity, including perception (Perc.), prediction (Pred.), planning (Pl.), control (C.) and end-to-end (E2E).
Figure 2: Overview of the trend of publishing datasets. The diagram exhibits a rapid growth of onboard datasets between 2015 and 2020 and then slowly down. In contrast, there is an increment trend for the V2X datasets, showing the research trend on the cooperative perception systems.
@@ I-B Main Contributions
  • We present an exhaustive survey on autonomous driving datasets. We consider publicly available datasets as comprehensively as possible, recording their fundamental characteristics, such as published year, data size, sensor modalities, sensing domains, geometrical and environmental conditions, and support tasks. To the best of our knowledge, we provide an overview of the most extensive collection of autonomous driving datasets recorded to date.
  • We systematically illustrate the sensors and sensing domains for collecting autonomous driving data. Furthermore, we describe the main tasks in autonomous driving, including task goals, required data modalities, and evaluation metrics.
  • We summarize and divide the datasets according to their sensing domains and supporting tasks, which assists researchers in efficiently selecting and gathering information for the goal datasets. Thereby facilitating more targeted and effective research and development efforts.
  • Additionally, we introduce an impact score metric to evaluate the influence of published perception datasets in the community. This metric can also be leveraged as a guidance for developing future datasets. We deeply analyze datasets with the highest grades based on the impact score, highlighting their advantages and utility.
@@
  • Detailed data statistics is conducted to demonstrate the data distribution of various datasets from different perspectives, exhibiting their inherent limitations and suitable use situations.
  • We analyze the recent technology trend and demonstrate the development direction of the next generation of datasets. We also provide a prospect on the potential effect of the advancement of Large Language Models on future autonomous driving.
@@ I-C Scope & Limitations
We aimed to conduct an exhaustive survey on the existing autonomous driving datasets to provide assistance and guidance for developing future algorithms and datasets in the domain. We collected datasets focusing on the four basic autonomous driving tasks: perception, prediction, planning, and control. Because several versatile datasets support multiple tasks, we only explained them in the main scope they support to avoid repeat introduction. Additionally, we collected a large number of datasets and exhibited them with their primary characters in tables. Nevertheless, a detailed explanation of all collected datasets could not highlight the most popular ones, and it can hurt assisting researchers in finding valuable datasets through this survey. Hence, we only describe the most impactful datasets in detail.
I-D Structure of the Article
The rest of the article is structured as follows: Section II introduces the approach we leveraged to obtain public datasets and the evaluation metrics for datasets. Section III demonstrates the primary sensors used in autonomous driving and their modalities. Section IV discusses autonomous driving tasks, related challenges, and required data. We further discuss several important datasets in Section V. The process of annotations and factors affecting annotation quality are exhibited in Section VI. Moreover, we statistic the data distribution of several datasets in section VII. In section VIII, we investigate the developing trend and future works of autonomous driving datasets. In the end, we conclude our work in Section IX. The taxonomy of this survey is shown in Fig. 3.
Figure 3: This survey’s primary taxonomy includes impact score, sensors and modalities, autonomous driving tasks, important datasets, and annotation process.
@@ II Methodology
This section consists of 1) How we collect and filter datasets (II-A), and 2) How to evaluate the impact of a dataset on the autonomous driving domain (II-B).
@@ II-A Datasets Collection
To ensure source diversity, we utilized well-known search engines such as Google [^1] https://www.google.com/, Google Scholar [^2] https://scholar.google.com/ and Baidu [^3] https://www.baidu.com/ to search datasets. To ensure a thorough dataset collection from various countries and regions, we conducted searches in English, Chinese, and German using keywords such as ”autonomous driving datasets,” ”intelligent vehicle datasets,” and terms related to object detection, classification, tracking, segmentation, prediction, planning, and control.
Furthermore, we explored IEEE Xplore [^4] https://ieeexplore.ieee.org/Xplore/home.jsp and pertinent conferences in autonomous driving and intelligent transportation systems to collate datasets from journals and conference proceedings. We verified datasets from these sources through keyword searches and manual title reviews.
Finally, to ensure the inclusion of specialized or lesser-known datasets, we searched through Github [^5] https://github.com/ repositories and Paperwithcodes [^6] https://paperswithcode.com/. Similar to databases, we performed both manual and keyword-based searches for datasets.
@@ II-B Dataset Evaluation Metrics
Citation Score. First, we calculate the citation scores from the total citation number and average annual citation. To gain fair citation counts, we choose the time of the earliest version of a dataset as its publish time. Moreover, all citation counts were collected as of September 20, 2023, to ensure the comparison is based on a consistent timeframe. The total citation number $\mathbf{c}^{t}$ reflects the overall influence of a dataset. A higher count of this metric means that the dataset has been widely recognized and utilized by researchers. However, datasets published in earlier years can accumulate more citations. To address this unfairness, we leverage average annual citation $\mathbf{c}^{a}$, which describes the yearly citation increase speed of a dataset. The calculate function is shown in Eq. 1.
$\mathbf{c}^{a}=\begin{cases}\mathbf{c}^{t}/\left(\mathbf{y}_{curr}-\mathbf{y}_{pub}\right)&\text{if }\mathbf{y}_{curr}\neq\mathbf{y}_{pub}\\ \mathbf{c}^{t}&\text{if }\mathbf{y}_{curr}=\mathbf{y}_{pub}\end{cases}$ (1)
@@
where $\mathbf{y}_{curr}$ and $\mathbf{y}_{pub}$ represent the current year and the dataset published year, respectively. On the other hand, the citation number of distastes has a wide distribution range from single digits to tens of thousands. To alleviate the extreme unbalance and highlight the differences of each dataset, we apply logarithmic transformation followed by Min-Max normalization to both $\mathbf{c}^{t}$ and $\mathbf{c}^{a}$, described in Eq. 2.
$\mathbf{c}_{norm}=\min\text{-}\max\left(\log(\mathbf{c})\right)$ (2)
Finally, the citation score $\mathbf{c}_{score}$ is the summation of $\mathbf{c}^{t}_{norm}$ and $\mathbf{c}^{a}_{norm}$:
$\mathbf{c}_{score}=0.5\*\mathbf{c}^{t}_{norm}+0.5\*\mathbf{c}^{a}_{norm}$ (3)
@@
Data Dimension Score. We measure the data dimension across four perspectives: dataset size, temporal information, task number, and labeled categories. Dataset size $\mathbf{f}$ is represented by the frame number of a dataset, reflecting its volume and comprehensiveness. To get the dataset size score $\mathbf{f_{normal}}$, we leverage the same method as the citation score to process the frame number to overcome the extreme imbalance between different datasets.
@@
Temporal information is essential for autonomous driving as it enables the vehicle to understand how the surrounding environment changes over time. We use $\mathbf{t}\in\left\{0,1\right\}$ to indicate whether a dataset includes temporal information. Regarding the task number, we only consider datasets related to the six fundamental tasks in the autonomous driving perception domain, such as 2D object detection, 3D object detection, 2D semantic segmentation, 3D semantic segmentation, tracking, and lane detection. Therefore, the task number score is recorded as $\mathbf{t_{n}}\in\left\{1,2,3,4,5,6\right\}$. A large number of categories is critical for the robustness and versatility of a dataset. During the statistic, if a dataset supports multiple tasks and includes various types of annotation, we choose the largest number of categories. Afterward, the categories are divided into five levels, $\mathbf{l}=\left\{1,2,3,4,5\right\}$, based on quintiles. To simplify the calculation, we normalize $\mathbf{t_{n}}$ and $\mathbf{l}$ before the following process.
@@
In order to reflect the data dimension score $\mathbf{d}_{score}$ as objectively as possible, we give different weights to the four components, as shown in Eq. 4.
$\mathbf{d}_{score}=0.5\*\mathbf{f}_{norm}+0.1\*\mathbf{t}+0.2\*\mathbf{t_{n}}_{norm}+0.2\*\mathbf{l}_{norm}$ (4)
@@
Environmental Diversity Score. We evaluate the environmental diversity of a dataset according to the following factors: 1) weather conditions, such as rain or snow. 2) time of day refers to when the data is collected, like morning or dusk. 3) the types of driving scenarios, e.g., urban or rural. 4) the geometric scope means the number of countries or cities where the data is recorded. It is worth noting that we treat the geometric scope for synthetic datasets as missing. We follow the granularity with which a paper categorizes its data to quantify the diversity. Moreover, for the missing value, if a dataset announces that the data is recorded under diversity conditions, we use the median value as the missing value. Otherwise, we set the missing value of this attribute to one. We apply quintiles to quantify each factor into five distinct levels. After that, the environmental diversity score $\mathbf{e}_{score}$ is the sum of these four factors.
In the end, we leverage Eq. 5 to calculate the impact score $\mathbf{i}_{score}$.
$\mathbf{i}_{score}=60\*\mathbf{c}_{score}+20\*\mathbf{d}_{score}+20\*\mathbf{e}_{score}$ (5)
@@
The total impact score is 100, and 60 percent belongs to the citation score $\mathbf{c}_{score}$, data dimension score $\mathbf{d}_{score}$ and environmental diversity score $\mathbf{e}_{score}$ takes 40 percent.
III Data sources and Cooperative Perception in Autonomous Driving

### v2

# A Survey on Autonomous Driving Datasets: Statistics, Annotation Quality, and a Future Outlook

@@ Abstract
Autonomous driving has rapidly developed and shown promising performance due to recent advances in hardware and deep learning techniques. High-quality datasets are fundamental for developing reliable autonomous driving algorithms. Previous dataset surveys either focused on a limited number or lacked detailed investigation of dataset characteristics. To this end, we present an exhaustive study of 265 autonomous driving datasets from multiple perspectives, including sensor modalities, data size, tasks, and contextual conditions. We introduce a novel metric to evaluate the impact of datasets, which can also be a guide for creating new datasets. Besides, we analyze the annotation processes, existing labeling tools, and the annotation quality of datasets, showing the importance of establishing a standard annotation pipeline. On the other hand, we thoroughly analyze the impact of geographical and adversarial environmental conditions on the performance of autonomous driving systems. Moreover, we exhibit the data distribution of several vital datasets and discuss their pros and cons accordingly. Finally, we discuss the current challenges and the development trend of the future autonomous driving datasets. https://github.com/MingyuLiu1/autonomous_driving_datasets.git
@@
Dataset, impact score, data analysis, annotation quality, autonomous driving.
@@ I Introduction
Autonomous driving (AD) aims to revolutionize the transportation system by creating vehicles that can accurately perceive their environment, make intelligent decisions, and drive safely without human intervention. Due to thrilling technical development, various autonomous driving products have been implemented in several fields, such as robotaxis [1]. These rapid advancements in autonomous driving rely heavily on extensive datasets, which help autonomous driving systems be robust and reliable in complex driving environments.
In recent years, there has been a significant increase in the quality and variety of autonomous driving datasets. The first apparent phenomenon in the development of datasets is the various data collection strategies, including synthetic datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] generated by simulators and recorded from the real world [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], to name just a few. Secondly, the datasets vary in composition, including but not limited to multiple sensor modalities like camera images and LiDAR point clouds, different annotation types for various tasks, and data distribution. Fig. 1 depicts the distribution of 3D object bounding box in six well-known real-world datasets (Argoverse 2[28], KITTI [13], nuScenes [22], ONCE [30], Waymo [23], and ZOD [31]) under a Bird’s-Eye View (BEV), highlighting each dataset’s unique annotation characteristics. The sensor mounting positions also reflect datasets’ various sensing domains, including onboard, Vehicle-to-Everything (V2X), or drone domain. The datasets’ geometric diversity and varying weather conditions also enhance the generalizability of autonomous driving datasets.
Figure 1: Bird’s-Eye View object distribution of datasets. Each heatmap represents a dataset and is plotted using X and Y coordinates. Y is the driving direction of the ego-vehicle. The unique annotation characters of each dataset are reflected in the distribution range, density, and number of bounding boxes.
@@ I-A Research Gap & Motivation
We demonstrate the yearly published number of perception datasets in Fig. 2 to illustrate the growth trends in autonomous driving datasets. Given the vast and growing number of publicly available datasets, a comprehensive survey of these resources is valuable for advancing both academic and industrial research in autonomous driving. In prior work, Yin et al. [32] summarized 27 publicly available datasets containing data collected on public roads. As a sequential work of  [32], [33] extended the number of datasets. Guo et al. [34] and Janai et al. [35] proposed systematical introductions to the existing datasets from an application perspective. Beyond describing existing datasets, Liu et al. [36] discussed the domain adaptation between synthetic and real data and automatic labeling methods. Li et al. [37] summarized existing datasets and undertook an exhaustive analysis of the characters of the next-generation datasets. However, these surveys only summarized a small number of datasets, causing a non-wide scope. AD-Dataset [38] collected a large number of datasets while lacking detailed analysis for the attributes of these datasets. Compared to studies on versatile datasets, some researchers presented surveys on a particular type of autonomous driving dataset, such as anomaly detection [39], synthetic datasets [40], 3D semantic segmentation [41], or decision-making [42]. Additionally, some task-specific surveys [43, 44] also organized the related AD datasets.
In this work, we present a comprehensive and systematic survey of a large number of datasets in autonomous driving. We compare our survey to others in Table I. Our survey covers all tasks from perception to control, considers real-world as well as synthetic data, and provides insights into the data modality and quality of several crucial datasets.
TABLE I: We compare our survey paper with other AD dataset surveys in the following perspectives: collected dataset number (#Dataset), relevant tasks, sensing domain (S. domain), sensor modality (S. moda.), geometric conditions (Geo.), environmental conditions (Env.), analyzing data distribution, introducing annotation quality and process. In the context of environmental conditions, we refer to the variability in weather conditions and illumination. Geometric conditions include scenario types and geographical scope. We describe the task types in a coarse granularity, including perception (Perc.), prediction (Pred.), planning (Pl.), control (C.), 0and end-to-end (E2E).
@@ A Survey on AD Datasets [36] 2021 30 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ours 2024 265 Perc, Pred, Pl, C, E2E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Figure 2: Overview of dataset publication trends from 2008 to 2024. The diagram demonstrates a significant increase in the publication of onboard datasets between 2015 and 2020, followed by a gradual decline thereafter. In contrast, there has been a rising trend in the publication of V2X datasets, indicating growing research interest in cooperative perception systems.
@@ I-B Main Contributions
  • We present an overview of the most exhaustive survey on autonomous driving datasets recorded to date. We show publicly available datasets as comprehensively as possible, recording their fundamental characteristics, such as published year, data size, sensor modalities, sensing domains, geometrical and environmental conditions, and support tasks.
  • We systematically illustrate the sensors and sensing domains for collecting AD data. Furthermore, we describe the main tasks in autonomous driving, including task goals, required data modalities, and evaluation metrics.
  • We categorize datasets based on their sensing domains and relevant tasks, enabling researchers to efficiently identify and compile information for their target datasets. This approach facilitates more focused and productive research and development efforts.
  • Additionally, we introduce an impact score metric to evaluate the influence of published perception datasets. This metric can also be a guide for developing future datasets. We deeply analyze datasets with high impact scores, highlighting their advantages and utility.
@@
  • Our detailed data statistics demonstrate the data distribution of various datasets from different perspectives, exhibiting their inherent limitations and suitable use cases.
  • We analyze recent technology trends and the development direction of next-generation datasets, such as integrating language into AD data, generating AD data using Vision Language Models, standardizing data creation, and promoting an open data ecosystem.
@@ I-C Scope & Limitations
We aim to conduct an exhaustive survey on the existing autonomous driving datasets to facilitate the development of future algorithms and datasets in this field. We collected datasets relevant to the five fundamental tasks of autonomous driving: perception, prediction, planning, control, and end-to-end (E2E) driving. To maintain clarity and prevent redundancy, we describe versatile datasets only within the main scope they support. Additionally, we collected a large number of datasets and exhibited them with their primary characters in tables. However, to ensure this survey aids researchers effectively, we focus our detailed discussions on the most impactful datasets rather than providing extensive descriptions of all datasets.
I-D Survey Structure
The rest of the survey is structured as follows: Section II introduces the approach leveraged to source public datasets and the evaluation metrics for them. Section III demonstrates the primary sensors used in autonomous driving and their modalities. Section IV discusses autonomous driving tasks, related challenges, and required data. In-depth discussions of several important datasets are presented in Section V. The process of annotations and factors affecting annotation quality are addressed in Section VI. Section VII provides statistical analysis of data distribution across various datasets. Future trends and potential research directions in autonomous driving datasets are explored in Section VIII. The paper concludes with Section IX. The survey’s taxonomy is shown in Fig. 3.
Figure 3: This survey’s primary taxonomy includes impact score, sensors and modalities, autonomous driving tasks, high-influence datasets, and annotation process.
@@ II Methodology
This section comprises two parts: 1) the collection and filtering of datasets, as detailed in Section II-A, and 2) the evaluation metric of a dataset’s impact on the autonomous driving domain, described in Section II-B.
@@ II-A Datasets Collection
To ensure source diversity, we utilized well-known search engines such as Google, Google Scholar and Baidu to search datasets. To ensure a thorough dataset collection from various countries and regions, we conducted searches in English, Chinese, and German using keywords such as “autonomous driving dataset/benchmark”, “intelligent vehicle dataset/benchmark” and terms related to object detection, classification, tracking, segmentation, prediction, planning, control, and end-to-end driving.
Furthermore, we explored IEEE Xplore, Paperswithcode, and pertinent conferences in autonomous driving and intelligent transportation systems to collate datasets from journals and conference proceedings. We verified datasets from these sources through keyword searches and manual title reviews.
Finally, to ensure the inclusion of specialized or lesser-known datasets, we searched through Github repositories. Similar to databases, we performed both manual and keyword-based searches for datasets.
@@ II-B Dataset Evaluation Metrics
Citation Score. First, we calculate the citation scores from the total citation number and average annual citation. To gain fair citation counts, we choose the time of the earliest version of a dataset as its publication time. Moreover, all citation counts were collected as of March 05, 2024, to ensure the comparison is based on a consistent timeframe. The total citation number $c^{t}$ reflects the overall influence of a dataset. A higher count of this metric means that the dataset has been widely recognized and utilized by researchers. However, datasets published in earlier years can accumulate more citations. To address this unfairness, we leverage average annual citation $c^{a}$, which describes a dataset’s yearly citation increase speed. The function is shown in Eq. 1.
$c^{a}=\begin{cases}c^{t}/\left(y_{curr}-y_{pub}\right)&\text{if }y_{curr}\neq y_{pub}\\ c^{t}&\text{if }y_{curr}=y_{pub}\end{cases}$ (1)
@@
where $y_{curr}$ and $y_{pub}$ represent the current year and the dataset published year, respectively. On the other hand, the citation number of distastes has a wide distribution range from single digits to tens of thousands. To alleviate the extreme imbalance and highlight the differences of each dataset, we apply a logarithmic transformation followed by Min-Max normalization to both $c^{t}$ and $c^{a}$, described in Eq. 2.
$c_{norm}=\min\text{-}\max\left(\log(c)\right)$ (2)
Finally, the citation score $c_{score}$ is the summation of $c^{t}_{norm}$ and $c^{a}_{norm}$:
$c_{score}=0.5\*c^{t}_{norm}+0.5\*c^{a}_{norm}$ (3)
@@
Data Dimension Score. We measure the data dimension across four perspectives: dataset size, temporal information, task number, and labeled categories. Dataset size $f$ is represented by the frame number of a dataset, reflecting its volume and comprehensiveness. To get the dataset size score $f_{norm}$, we leverage the same method as the citation score to process the frame number to overcome the extreme imbalance between different datasets.
@@
Temporal information is essential for autonomous driving as it enables the vehicle to understand how the surrounding environment changes over time. We use $t\in\left\{0,1\right\}$ to indicate whether a dataset includes temporal information. Regarding the task number, we only consider datasets related to the six fundamental tasks in the autonomous driving perception domain, such as 2D object detection, 3D object detection, 2D semantic segmentation, 3D semantic segmentation, tracking, and lane detection. Therefore, the task number score is recorded as ${t^{n}}\in\left\{1,2,3,4,5,6\right\}$. A large number of categories is critical for the robustness and versatility of a dataset. During the statistic, if a dataset supports multiple tasks and includes various types of annotation, we choose the largest number of categories. Afterward, the categories are divided into five levels, $cat=\left\{1,2,3,4,5\right\}$, based on quintiles. We normalize $t_{n}$ and $l$ before the following process to simplify the calculation.
@@
In order to reflect the data dimension score $d_{score}$ as objectively as possible, we give different weights to the four components, as shown in Eq. 4.
$d_{score}=0.5\*f_{norm}+0.1\*t+0.2\*t^{n}_{norm}+0.2\*cat_{norm}$ (4)
@@
Environmental Diversity Score. We evaluate the environmental diversity of a dataset according to the following factors: 1) weather conditions (such as rain or snow), 2) time of day (e.g., morning or dusk), 3) types of driving scenarios (like urban or rural), and 4) geometric scope refers to the number of countries or cities where the data is recorded. It is worth noting that we treat the geographical scope for synthetic datasets as undefined. We follow the granularity with which a paper categorizes its data to quantify the diversity. Moreover, for the missing value, if a dataset announces that the data is recorded under diversity conditions, we assign the median value; otherwise, the missing value is set to one. We apply quintiles to quantify each factor into five distinct levels. After that, the environmental diversity score $e_{score}$ is the sum of these four factors.
In the end, we leverage Eq. 5 to calculate the impact score $i_{score}$.
$i_{score}=60\*c_{score}+20\*d_{score}+20\*e_{score}$ (5)
@@
The total impact score is 100, and 60 percent belongs to the citation score $c_{score}$. Data dimension score $d_{score}$ and environmental diversity score $e_{score}$ takes 40 percent.
III Sensors and Perception Technology in Autonomous Driving